# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ 4.1 Latent factor
$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
Proposition 2.
Let everything be as above. Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\Delta_{N+1}=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}/\sum_{i\in\mathcal{M}}\nu_{i}.$
Proposition 2 shows that the presence of a common latent factor has little qualitative impact on the metric that captures the distortionary effects of misspecifed learning. The latter still equals the total misspecification associated with the sources in $\mathcal{M}$, weighted by their relative precision. Its comparative statics are the same as in the baseline model. In the leading example where the distortion of the manager’s worldview constitutes an unconscious negative bias against minorities, $\overline{\Delta}>0$. Proposition 2 predicts that in the long run, the DM undervalues minority workers’ productivity-enhancing investment by $\overline{\Delta}$, in addition to viewing sources favorable to them as too lenient and those adversarial toward them as impartial. Such a misperception clearly has impact on minority workers’ incentive to undertake costly investments, a subject we plan to explore in depth in a follow-up paper.
The proof of Proposition 2 shows that the transformed matrix $\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma$ is a rank-one update of a diagonal matrix, and so the KL minimization problem can be solved using an analogous approach to that developed in Section 3.3. Although it is challenging to exhaust all transformations of the DM’s problem that preserve the structural insights of the current analysis, it is certainly possible to consider specific transformations, such as one where the support of the DM’s prior (equivalently, the choice set in the KL minimization problem) is of the form ×$\times_{i\in\mathcal{S}}[\underline{b}_{i},\overline{b}_{i}]$. The baseline model is a special case where $\underline{b}_{i}=\overline{b}_{i}=\tilde{b}_{i}$ for $i\in\mathcal{M}$, whereas $\underline{b}_{i}=-\infty$ and $\overline{b}_{i}=+\infty$ for $i\in\mathcal{S}-\mathcal{M}$. It is easy to show, using the Lagrangian method, that this model variation has no qualitative impact on the distortionary effect of misspecified learning. All it does is endogenizing $\mathcal{M}$ as the set of sources whose interval constraints are binding; given $\mathcal{M}$, everything said so far remains valid.
@@ 4.2 Multidimensional state and signals
This section generalizes the baseline model to encompass multidimensional states and signals. In each period $t=1,2,\cdots,$ Nature draws the true state $\omega_{t}$ from a $K$-dimensional Gaussian distribution with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{0}$. Each dimension represents a distinct skill, such as problem solving, communication, teamwork, etc. The DM consults a set $\mathcal{S}=\{1,\cdots,N\}$ of information sources before making a prediction about the true state. The signal generated by source $i$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}^{K}$ is the $K$-dimensional bias of the source, and $\varepsilon_{i,t}$ is a source-specific, idiosyncratic error with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{i}$. The state and errors are jointly independent as in the baseline model, although different dimensions of the state or an error can exhibit correlations.
The true bias vector is $b^{*}\coloneqq[b_{1}^{*},\cdots,b_{N}^{*}]^{\top}\in\mathbb{R}^{NK}$. Initially, the DM misspecifies the biases of the sources in =$\mathcal{M}=\{1,\cdots,M\}$ as $\tilde{b}_{i}\neq b_{i}^{*}$, $i\in\mathcal{M}$, while being uncertain of the biases of the remaining sources. The support of his prior is $\{b\in\mathbb{R}^{NK}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. Everything else is as in the baseline model. In particular, the DM knows the true covariance matrix $\Sigma^{*}$ of the signals, which in the current setting equals $\text{diag}(\Omega_{1},\cdots,\Omega_{N})+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\otimes\Omega_{0}$ ($\otimes$ denotes the Kronecker product).
Key variables of our interest, including the distortion of DM’s long-run perceptions about source biases and that of his long-run behavior, are defined the same as before. The next proposition establishes the analog of Theorem 1 in the current multidimensional setting.
@@
The proof of Proposition 3 replaces the Sherman-Morrison formula with the Woodbury formula for inverting block matrices (Woodbury, 1950). Given the close resemblance between $\overline{\Delta}$ and its single-dimensional counterpart, we have no doubt that its comparative statics are generally ambiguous. However, we do know that $\overline{\Delta}$ continues to satisfy DOOM and so is independent of $\Omega_{i}$, $i\in\mathcal{S}-\mathcal{M}$. We could also gain insight into simple economic problems such as the one presented below, based on the closed-form formula for $\overline{\Delta}$.
@@
The DM faces a stream of female employees whose analytical skills and communication skills (hereafter denoted by $A$ and $B$, respectively) are drawn independently from the standard Normal distribution, i.e., $\Omega_{0}=\mathbb{I}_{2\times 2}$. The DM consults two sources about employees’ skills but misspecifies the biases of source one. Although the latter undervalues female employees’ analytical skills and overvalues their communication skills, the DM perceives it as impartial, i.e., $\Delta^{A}>0$ and $\Delta^{B}<0$. What the DM does know is that source one’s assessments of different skills exhibit a halo effect (Bohnet, 2016). That is, a positive impression about one’s analytical skill enhances the assessment of her communication skill, and vice versa. Let $\Omega_{1}=\small{\begin{bmatrix}v&v\rho\\ v\rho&v\end{bmatrix}}$ be the covariance matrix of the errors associated with source one’s signal. We capture the halo effect by letting $\rho>0$.
We examine how the DM’s misspecification about source one’s biases distorts his learning about source two’s biases, as well as his predictions about employees’ skills. Tedious algebra shows that
@@
In words, the DM ends up perceiving source two as too lenient about women’s analytical skills and too critical about their communication skills. This misperception is exacerbated by the halo effect, whose presence requires that the DM subtract the communication score from the analytical score to tease out the fundamentals from the errors. It is through this channel that the DM’s unconscious, negative, bias about the communication score assigned by source one creeps into his learning about the analytical score assigned by source two, making the latter appear more positively biased than it actually is.
@@
In the long run, the DM underestimates female employees’ analytical skills and overestimates their communication skills. To the extent that these estimates are key to HR decision making, such as task assignment, hiring, and retention, they can potentially reinforce occupational segregation and stereotypes that are pervasive in the workplace (see Blau and Kahn 2017 and Bertrand 2020 literature surveys). It is also natural for employees who anticipate such a misperception to overinvest in communication skills and to underinvest in analytical skills, thus fueling a vicious circle (Arrow, 1971). Our model uncovers the channels through which misspecified learning can potentially affect these labor market outcomes. A comprehensive exploration of these mechanisms is better left for future research.
@@
While it is tempting to suggest, based on the simple model presented above, that companies focus rater training on correcting for halo effects, complications may arise in the presence of multiple misspecified sources, as correcting for the halo effect exhibited by one source makes those exhibited by other sources more salient. The total effect could be an exacerbation or a mitigation of the distortions, for the same reason that renders the comparative statics of the signal-to-noise ratio ambiguous in Theorem 1(ii).
4.3 Unknown covariance matrix
So far we have assumed that the covariance matrix of the signals is known to the DM from the outset. In environments such as large organizations with numerous divisions and hierarchies, a more natural assumption seems to be that the covariance matrix is also a subject of misspecified learning. We entertain this possibility in this section, with a focus on whether the additional source of misspecified learning leads to more biased beliefs and behavior on the part of the DM, or a mitigation thereof.
Formally, fix the covariance matrix of Gaussian errors $\varepsilon_{1,t}\cdots,\varepsilon_{N,t}$ to any $\Omega^{*}\in\mathbb{R}^{N\times N}$ that is symmetric and positive definite (SPD). Fix any $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$, [^9] Throughout the paper, we use $a\wedge b$ to denote the minimum of $a$ and $b$ and $a\vee b$ to denote the maximum of $a$ and $b$. and define $\Lambda$ as the set of $N\times N$ SPD matrices whose minimum eigenvalues are at least $\underline{\lambda}$. Initially, the DM holds a prior belief about the distribution of the signals. The prior is correct about signals equal to their biases plus the true state and state-independent errors. However, it misspecifies the biases of some sources as in Section 2, while allowing the covariance matrix of the errors to be anything in $\Lambda$. The support of the prior, again denoted by $\Theta$, is
Importantly, $\Theta$ contains the true covariance matrix of the signals $\Sigma^{*}=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
We characterize outcomes of misspecified learning in two steps, beginning with the following lemma.

### v3

@@ 5.1 Unknown covariance matrix
[…]
@@ 5.2 Multidimensional state and signals
This section generalizes the baseline model to encompass multidimensional states and signals. In each period $t\in\mathbb{N}$, Nature draws the true state $\omega_{t}$ from a $K$-dimensional Gaussian distribution with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{0}$. Each dimension represents a distinct skill, such as problem solving, communication, teamwork, etc. The DM consults a set $\mathcal{S}=\{1,\cdots,N\}$ of information sources before making a prediction about the true state. The signal generated by source $i$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\epsilon_{i,t}$, where $b_{i}\in\mathbb{R}^{K}$ is the $K$-dimensional bias of the source, and $\epsilon_{i,t}$ is a state-independent Gaussian error with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{i}$. To focus on the correlations between different dimensions, we assume that errors are jointly independent across sources, while allowing different dimensions of an error and the true state to exhibit correlations.
The true bias vector is $b^{*}\coloneqq[b_{1}^{*},\cdots,b_{N}^{*}]^{\top}\in\mathbb{R}^{NK}$. Initially, the DM misspecifies the biases of the sources in =$\mathcal{M}=\{1,\cdots,M\}$ as $\tilde{b}_{i}\neq b_{i}^{*}$, $i\in\mathcal{M}$, while being uncertain of the biases of the remaining sources. The support of his prior is $\{b\in\mathbb{R}^{NK}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. Everything else is as in the baseline model. In particular, the DM knows the true covariance matrix of the signals, which in the current setting equals $\text{diag}(\Omega_{1},\cdots,\Omega_{N})+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\otimes\Omega_{0}$ ($\otimes$ denotes the Kronecker product).
The next proposition establishes the analog of Proposition 1 in the current multidimensional setting.
@@
The next example utilizes Proposition 3 to investigate the compound impact of misspecified learning and the halo effect in employee assessments.
@@
The DM faces a stream of female employees whose analytical and communication skills (hereinafter denoted by $A$ and $B$, respectively) are drawn independently from the standard normal distribution, i.e., $\Omega_{0}=\mathbb{I}_{2\times 2}$. The DM consults two sources about employees’ skills but misspecifies the biases of source one: Although the latter undervalues female employees’ analytical skill and overvalues their communication skill, the DM perceives it as impartial, i.e., $\Delta^{A}>0$ and $\Delta^{B}<0$. What the DM does know is that source one’s assessments of different skills exhibit a halo effect (Bohnet, 2016). That is, a positive assessment about one’s analytical skill enhances the assessment of her communication skill, and vice versa. Let $\Omega_{1}=\footnotesize{\begin{bmatrix}v&v\rho\\ v\rho&v\end{bmatrix}}$ be the covariance matrix of the errors associated with source one’s signal. We capture the halo effect by assuming that $\rho>0$.
We examine how the DM’s misspecification about source one distorts his learning about source two, as well as his prediction about employees’ skills, in the presence of the halo effect. Solving $\overline{\Delta}$ using Mathematica yields
@@
In words, the DM ends up perceiving source two as too lenient about women’s analytical skill and too critical about their communication skill. This misperception is exacerbated by the halo effect, whose presence requires that the DM subtract the communication score from the analytical score to discern the fundamental from the errors. It is through this channel that the DM’s unconscious, negative, bias about the communication score assigned by source one creeps into his learning about the analytical score assigned by source two, making the latter appear more positively biased than it actually is.
@@
In the long run, the DM underestimates female employees’ analytical skill and overestimates their communication skill. To the extent that these predictions are key to HR decision making, such as task assignment, hiring, and retention, they can potentially reinforce occupational segregation and stereotypes that are pervasive in the workplace (Bertrand, 2020). A comprehensive exploration of these mechanisms is better left for future research.
@@
While it is tempting to suggest, based on the simple model presented above, that companies focus rater training on correcting for halo effect, complications may arise in the presence of multiple misspecified sources, as correcting for the halo effect exhibited by one source makes those exhibited by other sources more salient. The total effect could be an exacerbation or a mitigation of the distortion, for the same reason that renders the comparative statics of signal precision ambiguous in Proposition 1.
6 Further applications
In the Online Appendix, we utilize our theoretical framework to analyze two additional applications: misperceived media bias and project investment under misspecified learning. In this section, we summarize our main findings, highlighting the implications of DOOM for policy interventions and regulations. Additional results and their analysis are relegated to the Online Appendix.
Misperceived media bias.
Consider a Republican voter learning about the horizontal bias of various news organizations. The misspecified sources consist of the voter’s personal sources and established networks. The former are viewed by the voter as unbiased due to bias blind spot (Pronin et al., 2002). The latter include, for instance, Fox News and CNN, which he incorrectly believes to be neutral, due to familiarity bias, and excessively left-leaning, due to liberal media bias, respectively (Pennycook and Rand, 2021; Hassell et al., 2020). These sources form the yardsticks with which he will learn about new entrants, including accounts on social media.
Our voter seeks to take a position that minimizes the quadratic distance from the truth. The optimal position in the long run is the BLP. In the Online Appendix, we fully describe the distortionary effects of misspecified learning on the voter’s long-run beliefs, behavior, and objective welfare (i.e., the expected quadratic loss under the true distribution). The welfare loss is $-\overline{\Delta}^{2}$ when signal variances are known to the DM, and $-\delta\overline{\Delta}^{2}$ when they are initially unknown. Importantly, both quantities satisfy DOOM.
DOOM implies that fact-checking interventions may have unintended welfare consequences. For one, sources with initially unknown biases are irrelevant to the distortionary effect of misspecified learning. By increasing the signal-to-noise ratio of these sources, the voter’s objective welfare increases unambiguously under both the correctly specified model and the misspecified model. [^14] The exact numbers are $-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}-\text{distortion}$, respectively. In contrast, increasing the signal-to-noise ratio of misspecified sources has ambiguous effects on $\overline{\Delta}^{2}$ and $\delta\overline{\Delta}^{2}$, as shown previously. Hence policy interventions that combat fake news by improving source accuracy may backfire when they modulate sources about which people already hold dogmatic misperceptions. Given that fact-checking is resource intensive, our model suggests prioritizing social media accounts over traditional media outlets, particularly when the specific nature of misspecifications is unclear to the fact-checker.
Project investment.
Consider a manager allocating capital across various projects. The manager can acquire information about project returns from various sources, ranging from internal analysts to uninvolved consultants and robo-advisors. Internal analysts, for instance, exhibit overoptimism bias that is neglected by the manager, due to the latter’s confirmation bias or the tendency to overly trust sources in his inner circle (Bieske et al., 2023; Malmendier, 2018). Uninvolved consultants and robo-advisors are novel and their biases are initially unknown.
We consider a manager with CARA utility and characterize the distortion in his portfolio choice and objective welfare due to misspecified learning. Interestingly, this is succinctly captured by terms depending on $\overline{\Delta}$. A consequence of DOOM is that the standard practice of seeking advice from “uninvolved consultants” as a means to counteract entrenched managerial biases (Guenzel and Malmendier, 2020) may offer limited benefits. Furthermore, technological advances that reduce data collection and processing costs could affect our manager differently, depending on whether they increase the signal-to-noise ratio of familiar or novel sources.
Appendix A Proofs
A.1 Preliminaries
Proof outline.
To begin, we describe the problem studied by HKS and discuss the relationship between our analyses.
HKS examine a misspecified Bayesian learning model in which the DM observes a $D$-dimensional Gaussian signal with mean $Mf\in\mathbb{R}^{D}$ and covariance matrix $\Sigma\in\mathbb{R}^{D\times D}$ in each period $t\in\mathbb{N}$. Initially, the DM knows the transformation matrix $M\in\mathbb{R}^{D\times N}$ — which captures group memberships — while misspecifying the first entry of the fundamental $f\in\mathbb{R}^{N}$. He may or may not know the true covariance matrix $\Sigma^{*}$. In Case (I) where $\Sigma^{*}$ is known, the support of the DM’s prior is
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1}\text{ and }\Sigma=\Sigma^{*}\}.$
In Case (III) where $\Sigma^{*}$ is initially unknown, the support becomes
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1}\text{ and }\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}}\},$
where $\underline{\underline{\lambda}}$ can be any positive number that is smaller than $1$ and $\lambda_{\min}(\Sigma^{*})$. Over time, the DM updates his beliefs about $(f,\Sigma)$ based on the prior and signal realizations. For both cases, HKS demonstrate that the DM’s limiting posterior beliefs concentrate on the KL minimizer and provide characterizations for the latter.
The main difference between our analysis and HKS is that our DM uses the outcome of misspecified learning to predict a random state. To facilitate analysis of the prediction problem, we make slightly different model assumptions from that of HKS, as outlined below:
  1. D1
    In our setting, $f=b$, $M=Identity$, and the DM can misspecify multiple entries of $b$.
  2. D2
    We assume that signals equal their biases, plus the true state and state-independent errors. The covariance matrix of the signals is thus a rank-one update of the covariance matrix of the errors, i.e., $\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. This structure is not imposed by HKS, who allow $\Sigma$ to be any SPD matrix with a bounded minimum eigenvalue.
How do these differences impact the analysis? For establishing the convergence of posterior beliefs to the KL minimizer, D1 proves to be inconsequential, whereas D2 requires that we bound the eigenvalues of $\Sigma$ based on the assumptions we make about $\Omega$. To this end, we apply spectral stability theories to bound the changes in the eigensystem that result from rank-one updates to an existing matrix. Details are in Appendix A.2, where we reproduce HKS’ convergence proof line by line.
When it comes to characterizing the limiting outcome, our proof can handle multi-sourced misspecifications and generalizes HKS’ guess-and-verify approach that is tailored for single-sourced misspecifications. When the covariance matrix is also a subject of learning, D2 introduces additional complications, as we need to convert the DM’s limiting belief about $\Sigma$ into a limiting belief about $\Omega$, using again spectral stability theories. Details are in Appendix A.3.
Notations.
We follow the notational conventions developed in the main text, using $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$ and $\text{diag}(x)$ to denote the diagonal matrix created from $x$. For each $k\in\mathbb{N}$, we use $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, and $\mathbb{I}_{k\times k}$ to denote the $k\times k$ identity matrix. Finally, $\otimes$ denotes the Kronecker product, $a\vee b$ stands for $\max\{a,b\}$, and $a\wedge b$ for $\min\{a,b\}$.
A.2 Useful lemmas and their proofs